# LLM-TabFlow: Synthetic Tabular Data Generation with Inter-column Logical Relationship Preservation

### v1

@@ Abstract
Synthetic tabular data have widespread applications in industrial domains such as healthcare, finance, and supply chains, owing to their potential to protect privacy and mitigate data scarcity. However, generating realistic synthetic tabular data while preserving inter-column logical relationships remains a significant challenge for the existing generative models. To address these challenges, we propose LLM-TabFlow, a novel approach that leverages Large Language Model (LLM) reasoning to capture complex inter-column relationships and compress tabular data, while using Score-based Diffusion to model the distribution of the compressed data in latent space. Additionally, we introduce an evaluation framework, which is absent in literature, to fairly assess the performance of synthetic tabular data generation methods in real-world contexts. Using this framework, we conduct extensive experiments on two real-world industrial datasets, evaluating LLM-TabFlow against other five baseline methods, including SMOTE (an interpolation-based approach) and other state-of-the-art generative models. Our results show that LLM-TabFlow outperforms all baselines, fully preserving inter-column relationships while achieving the best balance between data fidelity, utility, and privacy. This study is the first to explicitly address inter-column relationship preservation in synthetic tabular data generation, offering new insights for developing more realistic and reliable tabular data generation methods.
@@ 1 Introduction
In this paper, we thus aim to bridge this gap. First, we propose an evaluation framework to assess the effectiveness of synthetic tabular data in real-world scenarios. Next, we introduce LLM-TabFlow, a LLM-based synthetic tabular data generation method designed to preserve both privacy and inter-column relationships. This approach leverages LLMs to compress tabular data by inferring relationships and employs diffusion models to train and sample synthetic data in the latent space. Finally, we conduct experiments to comprehensively evaluate the proposed method and compare it with both traditional and state-of-the-art approaches. Experimental results on two real-world industrial datasets demonstrate that LLM-TabFlow outperforms existing methods under the proposed evaluation framework across three key dimensions: data fidelity, data utility, and data privacy, achieving a better trade-off between data fidelity and privacy preservation. By effectively preserving the complex inter-column dependencies, LLM-TabFlow is particularly well-suited for real-world applications beyond traditional machine learning tasks. It is the first method, to the best of our knowledge, to explicitly focus on capturing complex inter-column relationships in tabular data generation. By prioritizing relational consistency within tables, LLM-TabFlow sets a new benchmark for creating realistic and reliable synthetic datasets.
The rest of this paper is structured as follows. Section 2 reviews related work. Section 3 details the proposed approach, LLM-TabFlow. Section 4 presents experimental settings, including datasets, baselines, along with the evaluation framework. Section 5 presents experimental results. Section 6 discusses the limitation and implications of the proposed approach. Finally, Section 7 concludes the paper and summarizes future work.
@@ 2.2 Evaluation Framework
Currently, most research on tabular data generation methods is validated using low-dimensional datasets, such as Adult Income (Kohavi et al., 1996) and California Housing (Pace and Barry, 1997). In contrast, fewer studies explore large-scale datasets like HIGGS (Baldi et al., 2014) and Covertype (Asuncion et al., 2007). However, even these large-scale datasets predominantly consist of numerical attributes and lack complex inter-column relationships, particularly between categorical and numerical data. This limits their ability to represent real-world scenarios where logical consistency and dependencies are critical. And recent advancements in synthetic tabular data generation have been evaluated on real-world industrial datasets, particularly in finance and healthcare. For example, studies have used datasets such as Payments and Credit (Sattarov et al., 2023) and MIMIC-III (Johnson et al., 2016). These datasets inherently involve complex inter-column relationships, such as transaction amounts linked to account balances in financial data or patient diagnoses influencing treatment plans in healthcare. Despite the presence of these relationships, evaluations in these studies have largely overlooked their preservation. Instead, the focus has been on traditional metrics like machine learning utility and statistical alignment (Papadaki et al., 2024; van Breugel et al., 2024). Specifically, the current evaluation of data fidelity primarily focuses on capturing pairwise column correlations and multivariate dependencies (Margeloiu et al., 2024). While these metrics provide insights into statistical relationships, they fail to guarantee the consistency or preservation of logical relationships within multivariate distributions (Margeloiu et al., 2024). For example, in industrial tabular data from supply chains, complex inter-column relationships are often present, such as order dates being logically linked to delivery dates or total sales prices calculated as a function of price, quantity, and discounts (Suh et al., 2024; Borisov et al., 2022). Moreover, some logical relationships such as hierarchical consistency (e.g., city and country) are not unified. And conditional dependencies between those features may change depending on the specific combination of values in each row. This variability adds another layer of complexity to accurately modeling and evaluating these relationships. Recognizing this challenge, (Long et al., 2025b) proposed the first evaluation framework specifically designed for inter-column relationship preservation. Their work highlights the need for advanced generative models capable of accurately modeling these intricate, row-specific dependencies.
@@ 2.3 Tabular Data Generation Approaches
One of the earliest effective methods for generating tabular data is the Synthetic Minority Over-sampling TEchnique (SMOTE), originally proposed by Chawla et al. (2002) to address dataset imbalance. As a data augmentation method, this interpolation-based method has recently been found effective in generating tabular data (Wang et al., 2024a). Despite its simplicity, SMOTE achieves performance comparable to many advanced generative models on several benchmark datasets, as reported by Long et al. (2025b). However, since SMOTE generates new data points by linearly interpolating among neighboring data points without modeling the underlying data distribution, it inherently struggles to preserve data privacy effectively and is less capable of capturing complex relationships and dependencies between variables.
@@
Therefore, deep generative models have been introduced for tabular data generation. Xu et al. (2019) proposed TVAE, a Variational Autoencoder-based method, for generating tabular data. They further presented Conditional Generative Adversarial Networks (CTGAN), which achieved better performance than TVAE in terms of data accuracy. CTGAN transforms input categorical data into numerical formats using techniques such as one-hot encoding (Kim et al., 2022) and analog bits encoding (Zheng and Charoenphakdee, 2022). However, these encoding techniques often struggle to capture the rich dependencies and interactions in the input (Lee et al., 2023). This limitation restricts CTGAN from effectively recovering the complex logical relationships in real-world tabular data.
@@
Inspired by the remarkable success of diffusion models in in image (Ho et al., 2020) and audio generation (Huang et al., 2023), diffusion model-based methods have been proposed for tabular data generation, such as TabDDPM (Kotelnikov et al., 2023) and TabSyn (Zhang et al., 2023). Unlike TVAE and CTGAN, which rely on adversarial training and may struggle with issues like mode collapse or imbalanced data, diffusion model-based methods use a gradual denoising process, ensuring smoother and more stable learning of the tabular data distribution. These methods have shown to outperform TVAE and CTGAN in terms of better preserving inter-column relationships. However, the denoising process, an essential component of diffusion-model-based methods, is less effective for tabular data due to its heterogeneity, unlike image data, which is a more uniform in nature (Xu et al., 2019). Thus, diffusion models, with their isotropic noise structure, also struggles to effectively capture the logical relationships among columns (Long et al., 2025b).
More recently, the effectiveness of pretrained language models (PLMs) in natural language processing has led to the emergence of a new class of tabular data generation methods—LLM-based tabular data generation methods, such as GReaT (Borisov et al., 2022), REaLTabFormer (Solatorio and Dupriez, 2023), and Tabula (Zhao et al., 2023). These methods convert tabular data into tokenized formats before fine-tuning PLMs, enabling them to generate synthetic data that preserve the structure and semantics of the original dataset. However, compared to GAN-based and diffusion model-based methods, they require significantly longer fine-tuning and data generation times. Moreover, because PLMs rely heavily on sequential modeling, they are restricted their ability to model complex multivariate relationships when dependencies are not sequential. Therefore, the class of methods, GReaT, often fail to capture critical inter-column dependencies, reducing the fidelity of synthetic data. methods like GReaT often fail to capture critical inter-column dependencies, reducing the fidelity of synthetic data. While graph-based approaches, such as PAFT (Xu et al., 2024d), improve dependency preservation, they depend heavily on prior knowledge, which may not always be available or accurate.
@@
While existing methods demonstrate effectiveness in privacy preservation and in maintaining sequential inter-column relationships to some extent, they fall short in accurately preserving more complex inter-column dependencies in the generated synthetic data. We thus address this gap in the paper, proposing a LLM-based tabular data generation approach—LLM-TabFlow—which will be detailed in the following section.
3 The LLM-TabFlow Approach
@@
Recently, LLMs such as ChatGPT (Brown et al., 2020), LLaMa (Touvron et al., 2023), and Deepseek (Guo et al., 2025) have been demonstrated remarkable success in natural language processing and beyond. These models enable the automation of complex data processing tasks and significantly enhance the semantic understanding of text, including tabular data. For instance, Li et al. (2024) demonstrated their ability to summarize tabular data from reports, while Liu et al. (2023) showcased their potential for textual and symbolic reasoning on tables, facilitating deeper analysis and complex query resolution. These advancements highlight LLMs’ ability to bridge natural language understanding and structured data analysis. However, pretraining LLMs on large volumes of tabular text data (Hegselmann et al., 2023) demands significant computational resources and time. Additionally, tabular data lacks the structured context LLMs rely on, making them less effective for generation tasks compared to their strengths in text understanding. In contrast, deep generative models, such as diffusion models, excel at capturing complex patterns in tabular data by learning latent representations. These models efficiently model the joint distribution of features, enabling high-quality synthetic data generation. While deep generative models are adept at replicating statistical correlations, LLMs offer unique advantages in understanding semantic relationships and preserving inter-column dependencies.
Therefore, we propose LLM-TabFlow, which integrate LLMs reasoning into the deep generative models for tabular data generation. This method leverages deep generative models for latent space exploration while utilizing LLMs for semantic reasoning, ensuring both statistical accuray and domain-consistent relationships in the generated data. The workflow of LLM-TabFlow is illustrated in Figure 1, with each step detailed in the following sections.
Figure 1: Overview of the workflow of the LLM-TabFlow approach.
@@ 3.1 Serialization
The first step of LLM-TabFlow is serialization, as shown in Figure 1. Let $X=[x_{1},x_{2},\dots,x_{n}]$ represent the tabular dataset, where $x_{i}$ corresponds to the values of the $i$-th column, and $C=[c_{1},c_{2},\dots,c_{n}]$ denote the corresponding column names. Additionally, let $D=[d_{1},d_{2},\dots,d_{n}]$ denotes the column descriptions, where $d_{i}$ describes the semantics or metadata of the column $c_{i}$; $Y$ represents the target data (dependent variable). In this step, the column names and descriptions are serialized into natural language using the following mapping function:
@@ 3.2 LLM Reasoning
=$\mathcal{R}^{\mathcal{H}}=\Bigl{\{}c^{h}_{1,j}:\{c^{h}_{2,j},c^{h}_{3,j},\dots,c^{h}_{n,j}\}\Bigr{\}},$ (2)
@@
=$\mathcal{R}^{\mathcal{M}}=\Bigl{\{}\{f^{m}_{i_{1}},f^{m}_{i_{2}},\dots,f^{m}_{i_{n}}\}:\{f^{m}_{d_{1}},f^{m}_{d_{2}},\dots,f^{m}_{d_{q}}\}\Bigr{\}},$ (3)
@@ 3.3 Tabular Data Compression
The relationships $\mathcal{R}^{\mathcal{H}}$, $\mathcal{R}^{\mathcal{M}}$, and $\mathcal{R}^{\mathcal{T}}$ are preserved through the LLMs reasoning and integral to compressing complex tabular data, simplifying its structure and accelerating the subsequent data generation process. Specifically, $\mathcal{R}^{\mathcal{H}}$ is applied to identify and preserve columns at the most granular hierarchical level (e.g., “City”), while columns at higher levels in the hierarchy (e.g., “Region” and “Country”) are removed. Similarly, $\mathcal{R}^{\mathcal{M}}$ identifies independent columns within mathematical relationships and retains only those, discarding derived columns to reduce the complexity of the data structure. For temporal relationships, $\mathcal{R}^{\mathcal{T}}$ is used to retain the column with the earliest time point while transforming other columns within the temporal group $\{f^{t}_{1},f^{t}_{2},\dots,f^{2}_{n}\}$. The transformation involves computing the differences between consecutive time points, denoted as $dt_{k}=f^{t}_{k}-f^{t}_{k-1},\quad\text{for }k=2,3,\dots,p$, where $d^{t}_{k}$ represents the difference between the $k$-th time column and the $(k-1)$-th time column. This transformation simplifies the tabular data representation to be $\mathcal{X}_{\text{c}}$.

### v2

# LLM-TabLogic: Preserving Inter-Column Logical Relationships in Synthetic Tabular Data via Prompt-Guided Latent Diffusion

@@ Abstract
Synthetic tabular data are increasingly being used to replace real data, serving as an effective solution that simultaneously protects privacy and addresses data scarcity. However, in addition to preserving global statistical properties, synthetic datasets must also maintain domain-specific logical consistency—especially in complex systems like supply chains, where fields such as shipment dates, locations, and product categories must remain logically consistent for real-world usability. Existing generative models often overlook these inter-coloumn relationships, leading to unreliable synthetic tabular data in real-world application. To address these challenges, we propose LLM-TabLogic, a novel approach that leverages Large Language Model (LLM) reasoning to capture and compress the complex logical and relationships among tabular columns, while these conditional information are passed into a Score-based Diffusion for data generation in latent space. Through extensive experiments on real-world industrial datasets, we evaluate LLM-TabLogic for column reasoning and data generation, comparing it with five baselines including SMOTE and state-of-the-art generative models. Our results show that LLM-TabLogic demonstrates strong generalization in logical inference, achieving over 90% accuracy on unseen tables. And our method outperforms all baselines in data generation by fully preserves inter-column relationships while maintaining the best balance between data fidelity, utility, and privacy. This study presents the first method to effectively preserve inter-column relationships in synthetic tabular data generation without requiring domain knowledge, offering new insights for creating logically consistent real-world tabular data.
@@ 1 Introduction
In this paper, we thus aim to bridge this gap. First, we propose an evaluation framework to assess the effectiveness of synthetic tabular data in real-world scenarios. Next, we introduce LLM-TabLogic, a LLM-based synthetic tabular data generation method designed to preserve both privacy and inter-column relationships. This approach leverages LLMs to compress tabular data by inferring relationships and employs a score-based diffusion model to train and sample synthetic data in the latent space. Finally, we conduct experiments to comprehensively evaluate the proposed method and compare it with both traditional and state-of-the-art approaches. Experimental results on two real-world industrial datasets demonstrate that LLM-TabLogic outperforms existing methods under the proposed evaluation framework across three key dimensions: data fidelity, data utility, and data privacy, achieving a better trade-off between data fidelity and privacy preservation. By effectively preserving the complex inter-column dependencies, LLM-TabLogic is particularly well-suited for real-world applications beyond traditional machine learning tasks. It is the first method, to the best of our knowledge, to explicitly focus on capturing complex inter-column relationships in tabular data generation. By prioritizing relational consistency within tables, LLM-TabLogic sets a new benchmark for creating realistic and reliable synthetic datasets.
The rest of this paper is structured as follows. Section 2 reviews related work. Section 3 details the proposed approach, LLM-TabLogic. Section 4 presents experimental settings, including datasets, baselines, along with the evaluation framework. Section 7 presents experimental results. Section 8 discusses the limitation and implications of the proposed approach. Finally, Section 9 concludes the paper and summarizes future work.
@@ 2.2 Evaluation Framework
Currently, most research on tabular data generation methods is validated using low-dimensional datasets, such as Adult Income (Kohavi et al., 1996) and California Housing (Pace and Barry, 1997). In contrast, fewer studies explore large-scale datasets like HIGGS (Baldi et al., 2014) and Covertype (Asuncion et al., 2007). However, even these large-scale datasets predominantly consist of numerical attributes and lack complex inter-column relationships, particularly between categorical and numerical data. This limits their ability to represent real-world scenarios where logical consistency and dependencies are critical. And recent advancements in synthetic tabular data generation have been evaluated on real-world industrial datasets, particularly in finance and healthcare. For example, studies have used datasets such as Payments and Credit (Sattarov et al., 2023) and MIMIC-III (Johnson et al., 2016a). These datasets inherently involve complex inter-column relationships, such as transaction amounts linked to account balances in financial data or patient diagnoses influencing treatment plans in healthcare. Despite the presence of these relationships, evaluations in these studies have largely overlooked their preservation. Instead, the focus has been on traditional metrics like machine learning utility and statistical alignment (Papadaki et al., 2024; van Breugel et al., 2024). Specifically, the current evaluation of data fidelity primarily focuses on capturing pairwise column correlations and multivariate dependencies (Margeloiu et al., 2024). While these metrics provide insights into statistical relationships, they fail to guarantee the consistency or preservation of logical relationships within multivariate distributions (Margeloiu et al., 2024). For example, in industrial tabular data from supply chains, complex inter-column relationships are often present, such as order dates being logically linked to delivery dates or total sales prices calculated as a function of price, quantity, and discounts (Suh et al., 2024; Borisov et al., 2022). Moreover, some logical relationships such as hierarchical consistency (e.g., city and country) are not unified. And conditional dependencies between those features may change depending on the specific combination of values in each row. This variability adds another layer of complexity to accurately modeling and evaluating these relationships. Recognizing this challenge, (Long et al., 2025b) proposed the first evaluation framework specifically designed for inter-column relationship preservation. Their work highlights the need for advanced generative models capable of accurately modeling these intricate, row-specific dependencies.
@@ 2.3 Tabular Data Generation Approaches
One of the earliest effective methods for generating tabular data is the Synthetic Minority Over-sampling TEchnique (SMOTE), originally proposed by Chawla et al. (2002) to address dataset imbalance. As a data augmentation method, this interpolation-based method has recently been found effective in generating tabular data (Wang et al., 2024a). Despite its simplicity, SMOTE achieves performance comparable to many advanced generative models on several benchmark datasets, as reported by Long et al. (2025b). Mathematically, SMOTE can be defined as:
$\mathbf{x}_{new}=\mathbf{x}_{i}+\lambda(\mathbf{x}_{i}-\mathbf{x}_{j})$ (1)
@@
where $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are two randomly chosen instances from the minority class, and $\lambda$ is a random number between 0 and 1, controlling the distance between $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$. However, since SMOTE generates new data points by linearly interpolating among neighboring data points for each feature, it inherently struggles to capture complex relationships and dependencies among diverse columns. Therefore, deep generative models have been introduced for tabular data generation. Xu et al. (2019) proposed TVAE, a Variational Autoencoder-based method, for generating tabular data. They further presented Conditional Generative Adversarial Networks (CTGAN), which achieved better performance than TVAE in terms of data accuracy. CTGAN transforms input categorical data into numerical formats using techniques such as one-hot encoding (Kim et al., 2022) and analog bits encoding (Zheng and Charoenphakdee, 2022). The discriminator $D$ is trained to maximize the following objective:
$\mathcal{L}_{D}=-\mathbb{E}_{\mathbf{x}\sim p_{data}}[\log D(\mathbf{x})]-\mathbb{E}_{\mathbf{z}\sim p_{z}}[\log(1-D(G(\mathbf{z})))].$ (2)
@@
The discriminator $D$ in Eq. 2 is trained to distinguish between real data $\mathbf{x}$ and synthetic data $G(\mathbf{z})$, where $\mathbf{z}$ is drawn from a noise distribution $p_{z}$. This process focuses on optimizing the decision boundary between real and fake samples, capturing global patterns and correlations. However, it does not explicitly model discrete logical relationships, such as conditional dependencies (e.g., "if city = London, then country = UK"). These logical relationships require higher-level semantic understanding, which the discriminator, by focusing on broader correlations, cannot guarantee in the generated data. Inspired by the remarkable success of diffusion models in in image (Ho et al., 2020) and audio generation (Huang et al., 2023), diffusion model-based methods have been proposed for tabular data generation, such as TabDDPM (Kotelnikov et al., 2023) and TabSyn (Zhang et al., 2023). Unlike TVAE and CTGAN, which rely on adversarial training and may struggle with issues like mode collapse or imbalanced data, diffusion model-based methods use a gradual denoising process, ensuring smoother and more stable learning of the tabular data distribution. In the backward process (Equation 3), the transition from the original data $\mathbf{x}_{0}$ to the noisy data $\mathbf{x}_{t}$ is modeled as:
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (3)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ represents the cumulative noise scaling factor at step $t$. The backward process aims to denoise $\mathbf{x}_{t}$ to recover $\mathbf{x}_{0}$. For tabular data, even when embedded in a continuous space $\mathcal{X}$, feature relationships remain governed by discrete logical constraints rather than spatial correlations. Unlike natural images where neighboring pixels exhibit smooth variations (Xu et al., 2019), tabular embeddings may represent categorical variables or discontinuous relationships (e.g., $\text{age}<18\Rightarrow\text{is\_adult}=0$). This violates the local smoothness prior that makes diffusion models effective for images, requiring alternative approaches for tabular data denoising. Recent advances in pretrained language models (PLMs) have inspired LLM-based tabular data generation methods including GReaT (Borisov et al., 2022), REaLTabFormer (Solatorio and Dupriez, 2023), and Tabula (Zhao et al., 2023). These approaches tokenize tabular data and fine-tune PLMs using the objective function:
$L(\theta;T)=-\sum_{t_{i}\in T}\sum_{j=1}^{m}\log P(t_{i,j}|\Pi(t_{i},k)_{1:j-1}),$ (4)
where $P(t_{i,j}|\Pi(t_{i},k)_{1:j-1})$ represents the conditional probability of token prediction given preceding context. While this formulation effectively preserves inter-column logical relationships through autoregressive modeling, LLM-based methods face several challenges compared to GANs and diffusion models. The sequential nature of PLMs leads to substantially longer training and generation times on tabular data, while introducing sensitivity to feature ordering that can significantly impact model effectiveness (Xu et al., 2024d). Furthermore, the tokenization process and autoregressive objective struggle to accurately model continuous numerical distributions that frequently appear in tabular datasets, particularly for precise value generation in real-valued features.
@@
While existing methods demonstrate effectiveness in privacy preservation and in maintaining sequential inter-column relationships to some extent, they fall short in accurately preserving more complex inter-column dependencies in the generated synthetic data. We thus address this gap in the paper, proposing a LLM-based tabular data generation approach—LLM-TabLogic—which will be detailed in the following section.
3 LLM-TabLogic Approach
@@
Recently, LLMs such as ChatGPT (Brown et al., 2020), LLaMa (Touvron et al., 2023), and Deepseek (Guo et al., 2025) have been demonstrated remarkable success in natural language processing and beyond. These models enable the automation of complex data processing tasks and significantly enhance the semantic understanding of text, including tabular data. For instance, Li et al. (2024) demonstrated their ability to summarize tabular data from reports, while Liu et al. (2023) showcased their potential for textual and symbolic reasoning on tables, facilitating deeper analysis and complex query resolution. These advancements highlight LLMs’ ability to bridge natural language understanding and structured data analysis. However, pretraining LLMs on large volumes of tabular text data (Hegselmann et al., 2023) demands significant computational resources and time. Additionally, tabular data lacks the structured context LLMs rely on, making them less effective for generation tasks compared to their strengths in text understanding. In contrast, deep generative models, such as diffusion models, excel at capturing complex patterns in tabular data by learning latent representations. These models efficiently model the joint distribution of features, enabling high-quality synthetic data generation. While deep generative models are adept at replicating statistical correlations, LLMs offer unique advantages in understanding semantic relationships and preserving inter-column dependencies. Therefore, we propose LLM-TabLogic, a novel framework that addresses tabular data synthesis through two key innovations: First, we develop prompt-based methods to reason about and compress complex inter-column relationships while preserving their logical structure. Second, we demonstrate how these extracted conditional patterns can effectively guide generative models in the latent space to produce logically consistent synthetic data. The workflow of LLM-TabLogic is illustrated in Figure 1, with each step detailed in the following sections.
Figure 1: Overview of the workflow of the LLM-TabLogic approach.
@@ 3.1 Serialization
The first step of LLM-TabLogic is serialization, as shown in Figure 1. Let $X=[x_{1},x_{2},\dots,x_{n}]$ represent the tabular dataset, where $x_{i}$ corresponds to the values of the $i$-th column, and $C=[c_{1},c_{2},\dots,c_{n}]$ denote the corresponding column names. Additionally, let $D=[d_{1},d_{2},\dots,d_{n}]$ denotes the column descriptions, where $d_{i}$ describes the semantics or metadata of the column $c_{i}$; $Y$ represents the target data (dependent variable). In this step, the column names and descriptions are serialized into natural language using the following mapping function:
@@ 3.2 LLM Reasoning
$\mathcal{R}^{\mathcal{H}}=\Bigl{\{}c^{h}_{1,j}:\{c^{h}_{2,j},c^{h}_{3,j},\dots,c^{h}_{n,j}\}\Bigr{\}},$ (6)
@@
$\mathcal{R}^{\mathcal{M}}=\Bigl{\{}\{f^{m}_{i_{1}},f^{m}_{i_{2}},\dots,f^{m}_{i_{n}}\}:\{f^{m}_{d_{1}},f^{m}_{d_{2}},\dots,f^{m}_{d_{q}}\}\Bigr{\}},$ (7)
@@ 3.3 Tabular Data Compression
Many real-world datasets contain deterministic relationships—fixed rules where one variable is perfectly predictable from others, such as functional dependencies (City from ZIP Code), arithmetic constraints ($\textit{Total}=\sum\textit{Parts}$), or temporal sequences ($\textit{Departure\_Date}<\textit{Arrival\_Date}$). These relationships introduce no uncertainty ($P(X_{i}|X_{-i})=1$) and thus provide no meaningful signal for probabilistic models like diffusion or drift-based approaches to learn. Instead of wasting model capacity on reconstructing rigid, rule-based mappings, we can preprocess these dependencies out—either by removing existing derived variables or by replacing them with newly derived values. Recent advances have demonstrated large language models’ (LLMs) capabilities in tabular data processing, from summarization (Li et al., 2024) to symbolic reasoning (Liu et al., 2023). Building on these, we employ an LLM as a relational conditioner that compresses its logical relationships into a structured latent representation $\mathcal{R}$. This representation captures three fundamental dependencies: hierarchical ($\mathcal{R}^{\mathcal{H}}=\{c^{h}_{1}:\{c^{h}_{2},$...\}\}) through leaf nodes $X_{\ell}$ with invertible ancestor reconstruction $X_{a}=\phi_{a}(X_{\ell})$; mathematical ($\mathcal{R}^{\mathcal{M}}=\{\text{independent}\}:\{\text{derived}\}$) via minimal sufficient statistics $S=X_{j}$ and Dirac delta constraints $\delta(X_{k}-\psi(X_{j}))$; and temporal ($\mathcal{R}^{\mathcal{T}}=[f^{t}_{1},...,f^{t}_{n}]$) using initial states $X_{1}$ with Markovian increments $\{\Delta_{t}\}$. The reconstruction $\mathcal{D}=\mathcal{C}^{-1}$ preserves pairwise distributions $p_{\mathbf{X}}(x_{j},x_{k})$ through: (i) exact invertibility of $\phi_{a}$ and $\psi$ ensuring $p(\hat{x}_{j},\hat{x}_{k})=p(x_{j},x_{k})$, and (ii) Markov consistency via Chapman-Kolmogorov equations (contributors, 2023). And those Learned restoration functions $\{f_{decom}\}$ are subsequently used to condition our latent-space diffusion model during data generation.